OpenStudy (momobay212):

What is the fifth term of the geometric sequence? a1=120, a2=36, a3=10.8, a6=0.2916

hartnn (hartnn):

Hi @momobay212, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) Do you know how to find the common ratio?

OpenStudy (momobay212):

Thanks! And no I don't know how to do that, I basically just started doing geometric series.

hartnn (hartnn):

ok cool, then you might wanna have a look at this tutorial: http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0

hartnn (hartnn):

The common ratio is the ratio of 2nd term/ 1st term = 3rd term/ 2nd term = ... and so on

hartnn (hartnn):

here common ratio = r= a2/a1 = ...

OpenStudy (momobay212):

Ah ok, so I just divide the number that comes before the next?

hartnn (hartnn):

the next term, divided by the current term, yes

OpenStudy (momobay212):

How do I find the fifth term? Because there is a gap between the 3rd and 6th terms.

hartnn (hartnn):

so if we have the common ratio, we can get any term. multiple ways. we'll discuss one way here: r = a6/a5 we know a6, we found r

OpenStudy (momobay212):

0.3= 0.2916/a5

OpenStudy (momobay212):

is that right?

hartnn (hartnn):

r = a2/a1 = 36/120 = 0.3 correct. r = a6/a5 a5 = a6/r = 0.2916/0.3 = ...

OpenStudy (momobay212):

0.972

OpenStudy (momobay212):

Thank you so much!! :)

hartnn (hartnn):

correct! and welcome ^_^